# A Comprehensive Survey on Kolmogorov Arnold Networks (KAN)

### v3

@@ 6.1.1 Advantages of KAN in Symbolic Regression
The emergence of KAN has had a profound impact on the field of symbolic regression. Firstly, the high precision and interpretability of KAN allow researchers to better understand the mathematical relationships behind the data, thereby advancing scientific discovery. For example, Vaca-Rubio et al. (2024) demonstrated KAN’s superior performance in symbolic regression, showing through experiments that KAN is more accurate and efficient than traditional methods when handling complex datasets. Secondly, the flexibility and robustness of KAN make it particularly effective in dealing with noisy data. By dynamically adjusting activation functions, KAN can accurately identify the symbolic formulas that generate the data even in high-noise environments, providing higher reliability and stability for practical applications [5].
@@ 6.2.3 Experiments and Results
Research by Vaca-Rubio et al. (2024) also demonstrated the superior performance of KAN in satellite traffic prediction. Experiments showed that KAN models outperform traditional MLP models when dealing with complex satellite traffic data. KAN not only surpasses MLP in prediction accuracy but also demonstrates excellent parameter efficiency, achieving higher prediction performance with fewer trainable parameters [18].

### v4

@@ 6.1.1 Advantages of KAN in Symbolic Regression
KAN 2.0: Enhancing Symbolic Regression
Recent advancements in the field have further bolstered KAN’s capabilities in symbolic regression. Liu et al. (2024), in their study titled "KAN 2.0: Kolmogorov-Arnold Networks Meet Science," introduced several new tools such as MultKAN and the KAN compiler. These tools allow KAN to perform symbolic regression with even greater precision, especially in scientific contexts where uncovering underlying mathematical relationships is crucial. The dynamic adjustment capabilities provided by these tools enable KAN to better model and discover complex symbolic formulas, making it an invaluable asset in the realm of scientific discovery[19].
The emergence of KAN has had a profound impact on the field of symbolic regression. Firstly, the high precision and interpretability of KAN allow researchers to better understand the mathematical relationships behind the data, thereby advancing scientific discovery. For example, Kundu et al. (2024) demonstrated KAN’s superior performance in symbolic regression, showing through experiments that KAN is more accurate and efficient than traditional methods when handling complex datasets. Secondly, the flexibility and robustness of KAN make it particularly effective in dealing with noisy data. By dynamically adjusting activation functions, KAN can accurately identify the symbolic formulas that generate the data even in high-noise environments, providing higher reliability and stability for practical applications [5].
@@ 6.2.3 Experiments and Results
Research by Kundu et al. (2024) also demonstrated the superior performance of KAN in satellite traffic prediction. Experiments showed that KAN models outperform traditional MLP models when dealing with complex satellite traffic data. KAN not only surpasses MLP in prediction accuracy but also demonstrates excellent parameter efficiency, achieving higher prediction performance with fewer trainable parameters [18].
@@ 7.1 Review of Recent Research
KAN 2.0: Kolmogorov-Arnold Networks Meet Science
Recent research by Liu et al. (2024) has led to the development of KAN 2.0, a significant evolution of the Kolmogorov-Arnold Networks. This advancement focuses on integrating KAN with scientific methodologies, offering new functionalities such as MultKAN and the KAN compiler. These tools facilitate the extraction of scientific knowledge from data, making KAN more versatile in identifying relevant features, revealing modular structures, and discovering symbolic formulas[19].
the Robustness of Kolmogorov-Arnold Networks
In another critical development, Alter et al. (2024) explored the robustness of KAN in adversarial scenarios. Their work, titled "On the Robustness of Kolmogorov-Arnold Networks: An Adversarial Perspective," provides a thorough examination of KAN’s resilience against adversarial attacks, particularly in image classification tasks. This study highlights the potential of KAN as a secure and reliable model in environments where robustness is crucial, laying the groundwork for future research in enhancing KAN’s defensive capabilities[27].
@@ 8.3 Interdisciplinary Combinations and Research Opportunities
KAN 2.0 and Adversarial Robustness
Recent advancements like "KAN 2.0: Kolmogorov-Arnold Networks Meet Science" by Liu et al. (2024) provide new tools and methodologies that could significantly enhance KAN’s application in scientific discovery, particularly in symbolic regression. On the other hand, the study "On the Robustness of Kolmogorov-Arnold Networks: An Adversarial Perspective" by Alter et al. (2024) emphasizes the importance of further improving KAN’s robustness against adversarial attacks, which is critical for its application in security-sensitive areas such as finance and healthcare. Future research should focus on integrating these advancements to develop more robust, efficient, and versatile KAN models[27, 19].